OpenStudy (anonymous):

4-6). Solve the following systems by substitution. Substitute the expression in the second equation into the first equation and solve. 4x-y=5 y=-4x+1 what is x? what is y?

OpenStudy (anonymous):

4x+4x-1 = 5 8x=6 x=3/4 or 0.75 y=-2